# Exciton-Polariton Dynamics in Multilayered Materials

### v1

@@ Abstract
These three quantities, $\mathrm{Tr}[\hat{\rho}_{\mathrm{T}}^{2}(t)]$, $\mathrm{Tr}[\hat{\rho}_{\mathrm{EP}}^{2}(t)]$, and $\mathrm{Tr}[\hat{\rho}_{\mathrm{g}}^{2}(t)]=\rho_{\mathrm{gg}}^{2}(t)$, are presented in Fig. 4e-f at a cavity photon lifetime $t_{c}=263$ fs. Note that at $t\gg t_{c}$ the total purity reaches 1 because the total density matrix $\hat{\rho}_{\mathrm{T}}(t\gg t_{c})\rightarrow|\bar{0}\rangle\langle\bar{0}|$ which is a pure state. At the same time, the coherence in the exciton-polariton subspace becomes zero with $\hat{\rho}_{\mathrm{EP}}\rightarrow 0$. Overall the total purity (red solid line) decays till 250250~{}250250 fs primarily due to the decay of exciton polariton purity $\mathrm{Tr}[\hat{\rho}_{\mathrm{EP}}^{2}(t)]$. The rise in the total purity after $400$ fs is due to the increase in $\rho_{\mathrm{gg}}^{2}(t)$ as the population leaks to the ground state. Comparing to the no cavity loss scenario, represented by the violet and cyan solid line in Fig. 4e-f respectively, the $\mathrm{Tr}[\hat{\rho}_{\mathrm{EP}}^{2}(t)]$ decays faster, which is due to the loss in the total population in the exciton-polariton subspace. Importantly, regardless of the presence or absence of cavity loss, the multilayered material suppresses the decay in exciton-polariton purity.
This is further illustrated in Fig. 4i-j which presents the ratio of the purity in the multilayered to the single-layered scenario. We find that an increase in cavity loss suppresses the multilayer protection of coherence. This is expected as the phonon fluctuation synchronization effect in the multilayered materials only protects against phonon-induced decoherence and does not directly combat cavity loss. However, we find that even in a very lossy scenario ($t_{c}=131$ fs) multilayered materials can enhance purity by a factor of $\sim 4$.
Other parameter regimes. Overall, we find that the phonon fluctuation synchronization effect exists in a wide range of cavity setups. In the main text, we present two additional parameter regimes: (a) low photon frequency $\omega=720$ cm-1 and (b) low light-matter coupling regime with $\Omega_{0}=240$ meV.
Fig. 5 presents group velocity and exciton population density at $t=700$ fs (for the yellow-circled point in group velocity data) of single-layer and multilayered material (with $M$=25). In Fig. 5a-b, we use phonon frequency to $\omega=720$ cm-1 ($\gamma_{0}=1.1\sqrt{\omega^{3}}$). Just as the result presented above, Fig. 5a shows an enhancement in group velocity (over $\sim 50\%$) in multilayered materials compared to the single-layer case. Similarly, Fig. 5b shows that the exciton population density in multilayered materials exhibits a coherently propagating wavefront, which is absent in the single-layer case. In Fig. 5c-d, we set the exciton-photon coupling to $\Omega_{0}=240$ meV. For this parameter, Fig. 5c shows that the group velocity enhancement in multilayered materials, compared to single-layer materials, is small but still noticeable. However, Fig. 5d demonstrates that, for this exciton-photon coupling, exciton population density in the multilayer material propagates more coherently. The population density in the multilayered material are less noisy, and the population around the central site is more localized. The results for these two sets of parameters show that the coherence enhancement for the multilayered materials is a generic effect. However, having a larger exciton-photon coupling makes this enhancement more prominent.
Figure 5: Exciton-polariton dynamics in multilayered materials ($M=25$) at various parameter regimes. (a)-(b) exciton-polariton dynamics at photon frequency $\omega=720~{}\text{cm}^{-1}$ ($\gamma_{0}\approx 2.1\times 10^{-4}$ a.u.) with the rest of the parameters the same as Fig. 2d. (a) Exciton-polariton group velocity computed from quantum dynamical simulations vs. theoretical curve and (b) site-resolved exciton density at $t=700$ fs for yellow-circled point in (a). (c)-(d) Same as (a)-(b) but with light-matter coupling $\Omega_{0}=240~{}\text{meV}$ and $\omega=1440~{}\text{cm}^{-1}$ ($\gamma_{0}\approx 5.8\times 10^{-4}$ a.u.).
@@
Conclusion. In this work, we have studied the exciton-polariton dynamics in multilayer materials and compared their transport and coherence to single-layer materials. To achieve this, we introduced a new, fast, and highly efficient mixed quantum-classical approach that utilizes a bright layer description of the light-matter interactions, allowing us to simulate a large number of quantum states ($\sim 10^{6}$).
@@
Using our quantum dynamics approach, we found a phonon fluctuation synchronization effect where the delocalized light-matter interactions in a multilayered material shield exciton-polaritons from phonon-induced decoherence. As a result, exciton-polaritons propagate faster with an increased group velocity and, more importantly, remain more coherent for a longer time in multilayered materials despite interacting with a significantly larger number of phonon modes. We show that this phonon fluctuation synchronization effect originates from a fluctuation averaging in the transverse direction (perpendicular to the mirror plane) over multiple layers that suppresses dynamic disorder in the longitudinal direction (in the plane of the materials/mirrors) enhancing transport in the material plane. Our analytical analysis reveal that only one effective layer (the bright layer) couples with the cavity modes, while the rest remain uncoupled (dark layers). It is this bright layer, delocalized over all layers, that experiences lesser phonon fluctuations than in a single-layer material. We also show that even in the presence of cavity loss, multilayered material enhances transport and coherence, although to a lesser extent. We show such enhancement also exists in other parameter regimes, such as at lower phonon frequency or at lower light-matter coupling strength.
@@
Our work provides microscopic insights into exciton-polariton dynamics in multilayered materials and illustrates that quantum coherence lifetime can be extended by using multilayered cavity architechture. This work reveals that the pathway to polariton quantum devices is likely via stacked multilayered materials.
@@ III Acknowledgments
This work was supported by the Texas A&M startup funds. This work used TAMU FASTER at the Texas A&M University through allocation PHY230021 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. A.M. appreciates inspiring discussions with Yi Rao, Milan Delor, Daniel Tabor, Dong Hee Son, Pengfei Huo, David R. Reichman, Pritha Ghosh, and Sachith Wickramasinghe.

### v2

@@ Abstract
These three quantities, $\mathrm{Tr}[\hat{\rho}_{\mathrm{T}}^{2}(t)]$, $\mathrm{Tr}[\hat{\rho}_{\mathrm{g}}^{2}(t)]=\rho_{\mathrm{gg}}^{2}(t)$, and $\mathrm{Tr}[\hat{\rho}_{\mathrm{EP}}^{2}(t)]$, are presented in Fig. 4e-f at a cavity photon lifetime $t_{c}=0.53$ ps. Note that at $t\gg t_{c}$ the total purity reaches 1 because the total density matrix $\hat{\rho}_{\mathrm{T}}(t\gg t_{c})\rightarrow|\bar{0}\rangle\langle\bar{0}|$ which is a pure state. At the same time, the coherence in the exciton-polariton subspace becomes zero with $\hat{\rho}_{\mathrm{EP}}\rightarrow 0$. Overall, the total purity (red solid line) decays till $\approx 0.2$ ps primarily due to the decay of exciton polariton purity $\mathrm{Tr}[\hat{\rho}_{\mathrm{EP}}^{2}(t)]$. The rise in the total purity after $0.3$ ps is due to the increase in $\rho_{\mathrm{gg}}^{2}(t)$ as the population leaks to the ground state. Compared to the no cavity loss scenario, represented by the violet and cyan solid line in Fig. 4e-f respectively, the $\mathrm{Tr}[\hat{\rho}_{\mathrm{EP}}^{2}(t)]$ decays faster, which is due to the loss in the total population in the exciton-polariton subspace. Importantly, regardless of the presence or absence of cavity loss, the multilayered material suppresses the decay in exciton-polariton purity. This is expected as the phonon fluctuation synchronization effect in the multilayered materials only protects against phonon-induced decoherence and does not directly combat cavity loss.
Three-dimensional simulations. Here we demonstrate that the phonon fluctuation synchronization effect on the exciton-polariton dynamics is not limited to a simplified two dimensional light-matter system, and the same effect exists when considering the full three-dimensional light-matter Hamiltonian. For this, we consider two mirrors (DBR mirrors) placed in the $xy$-plane which are located at $z=0$ and $z=L$. For this cavity setup, we still consider the first cavity quantization mode along the $z$ direction such that $k_{z}=\pi/L$, therefore $\boldsymbol{k}=\boldsymbol{k}_{\parallel}+k_{z}\vec{e}_{z}=k_{x}\vec{e}_{x}+k_{y}\vec{e}_{y}+k_{z}\vec{e}_{z}$ for $\boldsymbol{k}_{\parallel}$ being the projection of the wavevector onto $xy$-plane. According to the solutions of Maxwell equations in three-dimensions for a Fabry-Pérot cavity, the cavity photons at each wavevector $\boldsymbol{k}$ consist of two cavity polarizations: the transverse electric field (TE or $s$) mode, and transverse magnetic field (TM or $p$) mode, which are degenerate for a bare cavity. For the bare exciton part of the light-matter Hamiltonian, we consider two-dimensional stacked multilayered material placed in parallel to the cavity mirrors. When considering a single-layer material, we place them at the middle of the optical cavity. For multilayered materials, additional layers are stacked above and below a layer placed in the middle of the optical cavity with an interlayer spacing of $a_{z}=4~\text{nm}$. The excitonic part of the system is described by a two-dimensional tight-binding model on a rectangular lattice with nearest-neighbor hopping. The lattice consists of $N_{x}$ sites along the $x$ direction with spacing $a_{x}$, and $N_{y}$ sites along the $y$ direction with spacing $a_{y}$.
Considering one exciton per unit cell, with a transition dipole aligned to the $y$ direction, the form of the light-matter interaction beyond the long-wavelength approximation is given by
$\begin{split}&\hat{H}_{\text{ex-cav}}=\\ &\frac{\Omega_{0}}{\sqrt{N}}\sum_{\bar{n},\boldsymbol{k}}\hat{X}^{\dagger}_{\bar{n}}\left(\sqrt{\frac{\omega_{\boldsymbol{k}}}{\omega_{0}}}{k_{y}}\hat{a}_{\boldsymbol{k}}-i\sqrt{\frac{\omega_{0}}{\omega_{\boldsymbol{k}}}}k_{x}\hat{b}_{\boldsymbol{k}}\right)\frac{\sin(k_{z}z_{m})}{|\boldsymbol{k}_{\parallel}|}e^{i\boldsymbol{k}_{\parallel}\cdot\boldsymbol{R}_{\bar{n}}}+h.c.\end{split}$ (18)
where $N=N_{x}N_{y}$ and $\hat{X}^{\dagger}_{\bar{n}}$ creates exciton at site $\boldsymbol{R}_{\bar{n}}=\boldsymbol{R}_{n_{x},n_{y},m}=n_{x}a_{x}\vec{e}_{x}+n_{y}a_{y}\vec{e}_{y}+ma_{z}\vec{e}_{z}$ for $n_{x}\in[0,N_{x}-1]$, $n_{y}\in[0,N_{y}-1]$, and $m$ is the layer index. Photon frequencies in this three-dimensional setup is given by $\omega_{\boldsymbol{k}}=\frac{c}{\eta}\sqrt{|\boldsymbol{k}_{\parallel}|^{2}+k_{z}^{2}}$ for $\eta=2.4$. Details of the light-matter Hamiltonian in a three-dimensional light-matter Hamiltonian are provided in the SI.
Fig. 5 presents the temporal exciton-polariton real-space population (density) in a three-dimensional optical cavity. For this figure, we use the exciton lattice constants of $a_{x}=a_{y}=20~\text{nm}$, $N_{x}=N_{y}=1000$, and all other parameters are the same as in Fig. 2-4. For the exciton-photon initial state, we model a resonant excitation in the lowest polariton band (defined by $\hat{P}_{k,0}^{\dagger}$ operator) at the energy $E_{0}$ by selecting an exciton-polariton wavefunction $|\Psi(0)\rangle=\sum_{k\in\mathcal{E}_{\mathrm{exc}}}\tilde{c}_{k}\hat{P}_{k,0}^{\dagger}|\bar{0}\rangle$, which is energetically localized within an energy window $\mathcal{E}_{\mathrm{exc}}$ (set to be $40$ meV around $E_{0}$). Details of the initial state and the split-operator approach for the full three-dimensional light-matter system is provided in the SI.
Fig.5a-c shows ballistic coherent propagation of exciton-polariton population in the absence of phonon ($\gamma=0$). Fig.5d-f demonstrates that for the single-layer material in the presence of exciton-phonon coupling ($\gamma=2\gamma_{0}$), exciton-polariton propagation is diffusive and incoherent, and the exciton-polariton population is localized near the center of the optical cavity where our initial state was prepared. However, the exciton-polaritonic dynamics in a multilayered material with same exciton-phonon coupling ($\gamma=2\gamma_{0}$) is ballistic and coherent as it is shown in Fig.5g-i. For a multilayer material, the propagation of the exciton-polariton population wavefront remains coherent beyond $t=0.658~\text{ps}$, which is presented in the SI. These results mirror the simpler two-dimensional scenario presented in Fig. 3, validating our understanding of phonon fluctuation synchronization effect in three-dimensions.
@@
Conclusion. In this work, we have studied the exciton-polariton dynamics in multilayer materials and compared their transport and coherence to single-layer materials. To achieve this, we introduced a new, fast, and highly efficient mixed quantum-classical approach that utilizes a bright layer description of the light-matter interactions, allowing us to simulate a large number of quantum states ($\approx 10^{10}$).
@@
Using our quantum dynamics approach, we found a phonon fluctuation synchronization effect where the delocalized light-matter interactions in a multilayered material shield exciton-polaritons from phonon-induced decoherence. As a result, exciton-polaritons propagate faster with an increased group velocity and, more importantly, remain more coherent for a longer time in multilayered materials despite interacting with a significantly larger number of phonon modes. We show that this phonon fluctuation synchronization effect originates from a fluctuation averaging in the transverse direction (perpendicular to the mirror plane) over multiple layers that suppresses dynamical disorder in the longitudinal direction (in the plane of the materials/mirrors), enhancing transport in the material plane. Mathematically we show that under an appropriate unitary transformation, only one effective layer (the bright layer) couples with the cavity modes, while the rest remain uncoupled (dark layers). Although, the phonons can transfer population between these dark layers and the bright layer or the cavity modes. It is this bright layer, delocalized over all layers, that experiences fewer phonon fluctuations than in a single-layer material. We also show that even in the presence of cavity loss, multilayered material enhances transport and coherence, although to a lesser extent.
@@
Our work provides microscopic insights into exciton-polariton dynamics in multilayered materials and illustrates that quantum coherence lifetime can be extended by using a multilayered cavity architecture. This work reveals that the pathway to polariton quantum devices is likely via stacked multilayered materials.
@@ III Acknowledgments
This work was supported by the Texas A&M startup funds. This work used TAMU FASTER at the Texas A&M University through allocations PHY230021 and PHY240260 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. A.M. appreciates inspiring discussions with Yi Rao, Milan Delor, Daniel Tabor, Dong Hee Son, Pengfei Huo, David R. Reichman, Pritha Ghosh, and Sachith Wickramasinghe.